# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

### v1

@@ 3.2 Embedding into a supergravity solution
$\mathrm{d}s^{2}$ $=R^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\mathrm{d}s^{2}_{T^{4}}+\mathrm{d}z^{2}$
$F$ $=2R^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z$
@@
$\mathrm{d}s^{2}$ $=R^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\bar{R}^{2}\left(\mathrm{d}y^{2}+\mathrm{d}z^{2}\right)+\mathrm{d}s^{2}_{T^{3}}$
$F$ $=2R^{2}\bar{R}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z.$ (3.4)
@@
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{\bar{R}}{R}\sigma^{0},0,r_{0}\right)$ (3.5)
@@
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is cured by the parameterisation of $t(\sigma)$ including the ration $\frac{\bar{R}}{R}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.13), one reproduces the MZW-model equations of motion for the fluctations $X_{S^{3}}^{m}$.
@@ 4 Outlook
Moreover, for the integrable two-dimensional $\sigma$-models, generalised geometry resp. O$(d,d)$-duality, was used in order to construct and understand new integrable models [41, 42, 43]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}/K_{d}$-structure, has been obtained and discussed in the literature [44, 45, 46, 47, 48, 49]. It might be fruitful to systematically study non-relativistic limits and post-Newtonian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [29].
@@ Acknowledgements
The author thanks the Institute for Theoretical Physics of ETH Zurich for their warm hospitality when finalising this work and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346.

### v2

@@ 3.2 Embedding into a supergravity solution
$\mathrm{d}s^{2}$ $=\bar{R}^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\mathrm{d}s^{2}_{T^{4}}+\mathrm{d}z^{2}$
$F$ $=2\bar{R}^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z$
@@
$\mathrm{d}s^{2}$ $=\bar{R}^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+R^{2}\left(\mathrm{d}y^{2}+\mathrm{d}z^{2}\right)+\mathrm{d}s^{2}_{T^{3}}$
$F$ $=2R\bar{R}^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z.$ (3.3)
@@
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{R}{\bar{R}}\sigma^{0},0,r_{0}\right)$ (3.4)
@@
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is cured by the parameterisation of $t(\sigma)$ including the ratio $\frac{R}{\bar{R}}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.7), one reproduces the MZW-model equations of motion for the fluctations $X_{S^{3}}^{m}$.
@@ 4 Outlook
Moreover, for the integrable two-dimensional $\sigma$-models, generalised geometry resp. O(d,d)(d,d)-duality, was used in order to construct and understand new integrable models [37, 38, 39]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}$-structure, has been obtained and discussed in the literature [40, 41, 42, 43, 44, 45]. It might be fruitful to systematically study non-Lorentzian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [30].
@@ Acknowledgements
The research of D.O. is part of project No. 2024/55/D/ST2/01205 funded by the National Science Centre (NCN) of Poland. The author thanks the Institute for Theoretical Physics of ETH Zurich for their warm hospitality when finalising this work, the participants of the ’Workshop on Higher-d Integrability’ in Favignana for fruitful discussions about the draft, and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346. Special credits belong to the anonymous referee of Physical Review Letters pointing out a crucial mistake in the original draft.
Appendix A The non-Lorentzian expansion of supergravity
The non-Lorentzian nature of the MZW-model suggests the use of non-Lorentzian geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [46, 47] and gives rise to the following expansions of metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter:
$g_{\mu\nu}$ $=c^{2}\tau_{\mu}{}^{A}\tau_{\nu}{}^{B}\eta_{AB}+c^{-1}H_{\mu\nu},$ (A.1)
$C_{\mu_{1}\mu_{2}\mu_{3}}$ $=c^{3}\tau_{\mu_{1}}{}^{A}\tau_{\mu_{2}}{}^{B}\tau_{\mu_{3}}{}^{C}\epsilon_{ABC}+C_{\mu_{1}\mu_{2}\mu_{3}}.$ (A.2)
Typically, the non-relativistic limit $c\rightarrow\infty$ was considered in the 11d supergravity case, but in other cases also the hyper-relativistic or Carollian limit ($c\rightarrow 0$) was considered from similar expansions in general relavitiy [48]. Both cases are of interest in main text.
$\eta_{AB}$ and $\epsilon_{ABC}$ are the $(1+2)$-dimensional Minkowski metric and totally antisymmetric tensor. The geometry, locally, describes a split into a $(1+2)$-dimensional relativistic and an eight-dimensional non-relativistic space, by three longitudinal (so-called clock) one-forms $\tau^{A}$ and a transverse metric $H$. In inverse of these objects is defined by the completeness and orthogonality conditions
$H^{\mu\rho}H_{\rho\nu}+{\tau^{\mu}}_{A}\tau_{\nu}{}^{A}$ $=\delta_{\nu}^{\mu},$
The relevant examples in this letter are very simplistic, allowing a global split into longitudinal and transverse space with coordinates $\hat{X}^{\mu}=(\hat{X}^{M},\hat{X}^{m})$. Moreover, $\mathrm{d}\tau^{A}=0$, the condition for the so-called Newton-Cartan torsionless membrane geometry, will always be satisfied here.